OpenStudy (anonymous):

Joshua is the place kicker for his college football team. Last season he kicked 42 times and never missed. Each field goal scored 3 points, and each extra point scored 1 point, for a total of 86 points. How many field goals did Joshua kick? A) 19 B) 20 C) 22 D) 23

OpenStudy (anonymous):

b

OpenStudy (anonymous):

hold on looking

OpenStudy (dtan5457):

f + p = 42 (Total number of kicks) 3f + p = 86 (Total number of points) Solve for p in the first equation p = 42 - f Now substitute that for p in the second equation 3f + (42 - f) = 86 Now just solve this equation for f 2f + 42 = 86 f = 22 Now we know f = 22, we can solve for p using either of the original equations 22 + p = 42 p = 20 Field goals = 22, Point after = 20

OpenStudy (anonymous):

its B

OpenStudy (dtan5457):

Nah he kicked 22.

OpenStudy (dtan5457):

It's C

OpenStudy (anonymous):

no @dtan5457 is right

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

so u mean b is not the right answer?

OpenStudy (dtan5457):

Nope

OpenStudy (anonymous):

did the math wrong sry :[

OpenStudy (dtan5457):

He had 20 extra points, but in reality he kicked 22.

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

@Esi7023 its C

OpenStudy (anonymous):

thxs yall:)

OpenStudy (anonymous):

now 3 more right

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok you can post on this page not a new one

OpenStudy (anonymous):

@Esi7023